A disoriented physics professor drives 3.25 km north, then 2.20 km west, and then 1.50 km south. Find the magnitude and direction of the resultant displacement using the component method.


1
Expert's answer
2022-01-10T09:15:47-0500

Find the magnitude and direction of the resultant displacement:


"D_N=3.25-1.5=1.75\text{ km}.\\
D_W=2.2\text{ km}.\\
D=\sqrt{D_N^2+D_W^2}=2.81\text{ km},\\\space\\
\theta=\arctan\dfrac{1.75}{2.2}=38.5°\text{ N of W.}"